OpenStudy (anonymous):

Please explain how L(1-cosφ) (line 2) changes to L[1-(1-(R/L)^2sin^2 θ)½] (line 3) 1) X = R + L – Rcos θ – Lsin φ 2) X = R (1-cosθ) + L(1-cosφ) 3) X = R (1-cosθ) + L[1-(1-(R/L)^2sin^2 θ)½] Where cosφ = (L2 – R2sin2θ)½/L = (1-(Rsinθ/L)^2)½ Kind regards

OpenStudy (anonymous):

it's just because they say that cosφ = (1-(Rsinθ/L)^2)½ \[\left( \frac{ Rsin(\theta) }{ L } \right)^2 = \frac{ R^2 }{ L^2 } \sin^2 \theta\]

OpenStudy (anonymous):

Thank you. So in other words (1-(Rsinθ/L)^2)½ is a trig identity of cosφ.

OpenStudy (anonymous):

I wouldn't call it a "trig" identity since phi and theta are undefined. But they do have a definite relationship

OpenStudy (anonymous):

Thanks Euler. Where do I look it up? Haven't seen it before.